OpenStudy (anonymous):

Use DeMoivre's Theorem to write the complex number in trigonometric form. (cos pi/4 + I sinpi/4)^3

zepdrix (zepdrix):

Ooo these are so fun :)

OpenStudy (anonymous):

I hope that means that you are good at them

zepdrix (zepdrix):

Hmm it's already in trig form... I guess we just apply DeMoivre's Theorem directly from here.\[\Large\rm \left(\cos \frac{\pi}{4}+\mathcal i \sin \frac{\pi}{4}\right)^3\]Applying DeMoivre, just bring the 3 inside,\[\Large\rm \left(\cos \frac{3\pi}{4}+\mathcal i \sin \frac{3\pi}{4}\right)\]Yah?

OpenStudy (anonymous):

well you just made it look easy thanks

zepdrix (zepdrix):

They get a little trickier when they give them to you in a+bi form and make you convert to trig before you can apply DeMoivre... But yah they start out nice and easy! :)

OpenStudy (anonymous):

\[\left( \cos \theta +\iota \sin \theta \right)^n=\cos n \theta+\iota \sin n \theta\]